OpenStudy (dschneider2016):

Please help!! Will medal!! Verify the identity. cot(x- pi/2) = -tanx

OpenStudy (dschneider2016):

@priyar @nincompoop @dramirez9 Please help!

OpenStudy (priyar):

do u know cot(pi/2 - x) =tanx ?

OpenStudy (dschneider2016):

no does that mean the same thing?

OpenStudy (priyar):

not exactly the same thing..but u must have learnt this right?

OpenStudy (dschneider2016):

I think I have seen it before

OpenStudy (priyar):

ok..and do u know cot(-x) = -cotx ?

OpenStudy (dschneider2016):

yes, so what is the first step for this problem

OpenStudy (priyar):

these are the only things u need to solve this problem!

OpenStudy (priyar):

cot(pi/2 - x) ..take -1 common out from inside the bracket..what do u get?

OpenStudy (priyar):

@dschneider2016 ?

OpenStudy (dschneider2016):

umm let me see, -cot(pi/2 - x) =tanx instead of cot(pi/2 - x)= -tanx ? @priyar

OpenStudy (priyar):

no..rewrite cot(pi/2 - x) after taking -1 common out

OpenStudy (priyar):

let-1 be just outside the brackets...

OpenStudy (dschneider2016):

do you mean -cot(pi/2 - x)

OpenStudy (priyar):

no.i mean cot-(x - pi/2)

OpenStudy (dschneider2016):

oh that makes sense. What is the next step @priyar

OpenStudy (priyar):

now.. this is in the form of cot( -y )..so what can this be written as (see abv.)

OpenStudy (dschneider2016):

-cotx

OpenStudy (priyar):

is it "x" over there?

OpenStudy (dschneider2016):

no so it would be -coty?

OpenStudy (priyar):

here what is our y?

OpenStudy (priyar):

so cot-(x - pi/2) can be written as -cot(x - pi/2) do u agree?

OpenStudy (dschneider2016):

y= (x - pi/2)

OpenStudy (dschneider2016):

yes I do

OpenStudy (priyar):

good

OpenStudy (priyar):

no.. i made a mistake ..

OpenStudy (dschneider2016):

oh what as it?

OpenStudy (priyar):

so -cot(pi/2-x)=cot(x- pi/2) do u agree? i meant to ask this..

OpenStudy (dschneider2016):

yes I do

OpenStudy (priyar):

did u understand how this came we have just taken the negative sign outside..

OpenStudy (dschneider2016):

yes but is this all we have to do to verify the identity?

OpenStudy (priyar):

and we already know cot(pi/2 - x) = tanx put that in the abv expression..

OpenStudy (priyar):

i mean this exp : -cot(pi/2-x)=cot(x- pi/2)

OpenStudy (dschneider2016):

Im kinda confused. If we were to combine everything to create a clear justification for verifing this identity what would it be (all the way to the answer)? @priyar @dumbcow

OpenStudy (priyar):

ok let me make things clearer..wait..

OpenStudy (priyar):

We know , cot(pi/2-x)=tanx => -cot(x-pi/2)=tanx => cot(x-pi/2) = -tanx

OpenStudy (priyar):

hence verified

OpenStudy (dschneider2016):

so thats all I need?

OpenStudy (priyar):

yes..the explantion behind how these steps came is what i was making u understand..and u did pretty well..so i hope u understood..how to start when u see a problem like this and the 2 identities that we have used here..

OpenStudy (dschneider2016):

yes I understand now! Thank you so much! you have been literally so helpful!

OpenStudy (priyar):

u r welcome! be sure to learn all the trig identities and remember them you will be using a lot of them frequently in maths...

OpenStudy (dschneider2016):

yes! thanks so much! Im going to try to memorize them now

OpenStudy (priyar):

thats good!

OpenStudy (whpalmer4):

\[ \cot(x- \pi/2) = -\tan x\]break down into component functions \[\frac{\cos(x-\pi/2)}{\sin(x-\pi/2)} = -\frac{\sin x}{\cos x}\]but \(\sin x\) lags \(\cos x\) by \(\pi/2\) so \(\cos(x-\pi/2) = \sin x\) \[\frac{\sin x}{\sin(x-\pi/2)} = -\frac{\sin x}{\cos x}\] similarly, \(\sin(x-\pi/2) = -\cos x\) so \[\frac{\sin x}{-\cos x} = -\frac{\sin x}{\cos x}\]which is clearly true, so the identity is verified.